OpenStudy (kobeni-chan):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4> a. 20.3 b. 10.2 c. 0.2 d. 30.3

OpenStudy (kobeni-chan):

I know that I use \[\cos \theta = \frac{ u \times v }{ \left| \left| u \right| \right| \left| \left| v \right| \right| }\] and right now I have \[\frac{ 42, 4 }{ \sqrt{37} \sqrt{65} }\]

OpenStudy (kobeni-chan):

@mathmale

OpenStudy (kobeni-chan):

@inkyvoyd @johnweldon1993